A Bibliography of Tree Drawing Algorithms

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 3.28

Abstract

This bibliography collects papers and reports on algorithms for the drawing of trees.

Title word cross-reference

4 [KH97]. h [MPS94]. k [Val97]. O(n^{2}) [CS97]. v [MPS94].

-connected [KH97].

1st [NS87].

2 [FKK97]. 2- [CS97]. 2-Visibility [FKK97]. 2D [ARRM97]. 2nd [ACM89].

3 [BSWW97, DLV93, LD95]. 3- [CS97]. 3-D [BSWW97]. 3-Dimensional [LD95].

3-Planar [DLV93]. 3D [ARRM97, CT95]. 3DCube [PV97].

4-colorable [CS97].

5th [Di 97].

’87 [NS87].

’91 [MK91]. ’94 [TT95]. ’95 [Bra96]. ’96 [Nor97].

action [KR91]. Acyclic [CG95, DGLP97, DT88, HL96]. Ada [BL86]. Adocs [Ber95]. Advances [GT94a]. Aesthetics [PCJ95]. Agraph [Nor02a]. aided [Car80]. Alberta [MK91]. Algorithm [BFGT97, CP95a, GT97, KT97, KRB95, RNN97, See97, SM95b, Thi96, CP95b, KK89, MAUY98]. Algorithms [BHR95, BDLT95, DT88, EFL97, ESW97, FK97a, Foe97, GT94b, PT95a, SSH^{+}98, SKC96, SKC00, Til81]. analysis [Car80]. Angle [Gar95]. Angles [DV96]. Angular [GT94b]. Animated [ECH97]. Annealing [CT95, DH96]. Annual [ACM90]. Applications [Cla90, GT97, NK94, GN00, KH97]. Approach [BMT97, MR95, MB91]. approximability [Pen02]. Arbitrary [GR02]. Archaeology [HMPT97]. ArchE [HMPT97]. Area [BK97, CGKT97, Cha02, CP95c, CP97a, FK97a, GR02, LE97, PT98, SKC96, SKC00, Kim97, Tre96b, Tre96a]. Area-Efficient [SKC96, BK97, SKC00]. ary [Rad88]. Aspect [CGKT97, DGK98, GR02]. Attribute [ZM96]. Automatic [Car80, JLM97, Par95, See97, TdB88]. Autostereograms [Thi96]. AVL [CP95c, Kim97].

Balanced [DGK98]. Barzdin [ESW96]. Based [ARRM97, NT87, SM95a]. Bend [FK95b, FK95a]. Bend-Minimum [FK95b]. Bends [BDD98, DL97, PT98]. Berkeley [Nor97]. Best [EHW97]. Better [BK94]. Binary [Cha02, CP97a, GR02, MPS94, SSH^{+}98]. Bipartite [FK97b]. books [MB91]. Bound [Cha02]. Bounds [Bie95, FK97a, Foe97, GT94b, PT95a, TTV91]. Browser [RDM^{+}87]. Browsers [ARRM97].

Calgary [MK91]. California [Nor97]. CG [MSG95]. charts [BL86]. chemical [Fuj95]. circuit [MB91]. Clan [SM95a]. Clan-Based [SM95a]. Class [See97, Thi96]. Clustered [EF97, EFL97, FCE95]. Clustering [New89]. colorable [CS97]. Combinatorial [Ber95]. Comparison [BHR95]. Complete [CP97a, Tre96b, Tre96a]. Complexity [SR83, ABJT04]. Compound [Par95]. computation [GS85]. computer [Car80, KR91]. computer-supported [KR91]. Computing [BDD98, DL97, LD95]. Concentration [New89]. Concept [SW95]. Conference [NS87]. Configuration [ACM89]. Configurations [Gro95]. connected [KH97]. construction [VD93, VD94]. Contest [EM95]. conventions [Pen02]. Convexly [Dji95]. cooperative [KR91]. corrigendum [ABJT04]. Cost [GT97]. Crossing [Val97, War77]. Cubic [CP95a, RNN97].

D [BSWW97]. DAG [GNV88]. Data [Mal82, Fra81]. decision [Car80]. Decomposition [SM95a]. Definition [vOV96]. Degree [FK95a, PT98]. Depth [MAUY98]. Depth-first [MAUY98]. Design [ARRM97, KR91]. Designing [Bra95]. Diag [Sey92]. Diagrams [Lan82, Lev90, See97, Thi97, GS85, MB91, Ohl95, TdB88]. Diagraphs [CG95]. Digraphs [CDTT95, Par95, DT88, HL96, CDTT95]. DIMACS [TT95]. Dimensional [CELR95, EHW97, ESW97, PV97, DN97, ESW96, PTT97, LD95]. Dimensions [EG95, HE03, PT97]. Directed [BHR95, DGLP97, FK95b, GKNV93, Ign95, NT87, New89, RDM^{+}87, SM95a, Fra81, FR91, GNV88]. display [Car80]. dot [KN91, GKN02, Kou96a]. dotty [Kou94a, Kou96b]. Draw [FCE95, LL95]. Drawable [ELLM95, LL97]. drawer [Swo91]. Drawing [ARRM97, BDGS97, Ber95, BSWW97, BMT97, BGRT95, BHR95, BGT97, BFGT97, BKW89, BKW90, BJL06, BL86, BDLT95, CP95a, CS97, Cha02, CELR95, CG95, CT95, DH96, DN97, DGLP97, Dji95, DGK98, Ead84, EW94, EG95, EM95, Ead96, EF97, ECH97, EFL97, ESW97, FK95b, FK95a, Foe97, FHH^{+}93, FR91, GKNV93, GKN02, Gan03, GT94a, Gar95, GT97, Gro95, HIIN99, HE03, Hop71, HMPT97, Ign95, JRZ95, JLM97, KRB95, KN91, Kou96a, Lan82, Leg91, LE97, MACL95, MSG95, Moe90, Nor02b, Ohl95, PT95b, PT97, PT98, Par95, PV97, PCJ95, Rad88, RT81, Saa95, See97, SM95a, SW95, SM95b, SR83, Thi97, Thi96, TX95, TBC97, WS79, ZM96, Žit94, vdP14, ABJT04, Bra96, CJKC00, CP95b, Di 97, Fuj95, HL96, KK89, KH97]. drawing [Kim97, Lev90, MB91, Nor97, Pen02, SSH^{+}98, Sey92, TdB88, TT95, Til81, Tol96, Tre96b, Tre96a, VD93, VD94]. Drawings [BDD98, BK94, Bie95, Bie96, Bra95, CGKT97, CN95, CDTT95, CP95c, CP97a, CP97b, DLV93, DL97, EHW97, FK97a, FKK97, FK97b, GT94b, GR02, HKR95, HK95, KT97, LD95, MPS94, MR95, PTT97, PT95a, RNN96, RNN97, SKC96, Val97, BK97, ESW96, SKC00, TTV91, VA88]. draws [GNV88]. DynaDAG [Norxx]. Dynamic [CDTT95, EN96, Moe90].

Edge [HK95, New89, KH97]. Edges [KT97, Val97]. Editing [Kou94a, Kou94b, Kou96b, Kou96c, Fra81, Lew75]. editor [KS90, KD91]. Editors [NT87]. Efficient [CP95a, SKC96, BK97, SKC00]. Efficiently [CG95]. Empirical [Swe78]. encoding [DN97]. Engineering [NS87, GN00]. entropy [Swe78]. environment [VD94]. Equivalence [Thi96]. ESEC [NS87]. estimates [Swe78]. European [NS87]. event [KR91]. event-action [KR91]. evolutionary [VD93, VD94]. Example [CG95]. Experimental [BHR95, DGLP97]. Extending [See97]. extensible [KS90]. Extension [EN96].

Few [PT98]. Feynman [Lev90, Ohl95]. Fibonacci [Tre96b, Tre96a]. Finding [EHW97]. first [MAUY98]. Flow [GT97, MR95]. Forbidden [LL97]. Force [BHR95, FR91, Ign95]. Force-Directed [BHR95, Ign95, FR91]. Formal [Mal82]. formulas [Fuj95]. France [NS87].

Gallery [ARRM97]. GD

[Bra96, BDLT95, Nor97, TT95].

GD-Workbench [BDLT95]. GD’97 [Di 97]. General [Wal80, Wal91, Fra81, GS85, KK89]. Generalization [Saa95]. Generic [Ber95]. Geometric [MACL95]. Geometrical [SW95]. Germany [Bra96]. Grammar [MSG95]. Grammars [Bra95, ZM96]. Graph [ARRM97, BDGS97, BK94, Bie95, BSWW97, BMT97, Bra95, BHR95, Bra96, BGT97, BFGT97, BDLT95, CDTT95, CELR95, CT95, Di 97, Dji95, Ead84, EM95, Ead96, EHW97, ECH97, ESW97, FCE95, Foe97, FR91, GT94a, GT97, Hop71, HMPT97, JLM97, MSG95, NK94, Nor97, PT95b, PT97, PV97, PCJ95, Saa95, TT95, Tol96, TBC97, Val97, ZM96, BK97, CP95b, ESW96, GN00, KH97, KS90, TdB88]. Graph-Grammar [MSG95]. Graphed [Him9x]. Graphic [Mal82]. Graphics [Lev90, MK91]. Graphs [Bie96, CP95a, CS97, CN95, CG95, DLV93, DV96, DGLP97, DGK98, EG95, EF97, EFL97, ELLM95, EN96, FK95b, FK95a, FKK97, FK97b, FHH^{+}93, GKNV93, Gar95, HK95, Ign95, KRB95, LE97, MACL95, NT87, New89, PTT97, PT98, RNN96, RNN97, RDM^{+}87, SM95a, SM95b, ZM96, Žit94, Car80, DH96, EW94, GNV88, GKN02, Gan03, KK89, KH97, KN91, Kou94a, Kou96a, Kou96b, Nor02b, TTV91]. GraphViz [Gan03]. Grid [CP95a, CN95, EF97, JRZ95, PTT97, RNN96, CP95b, DL97]. Guide [Nor92]. GWEZ [Leg91].

h [CP97a]. h-v [CP97a]. Haskell [Thi97]. Heuristic [BK94, Ead84, SW95]. Hierarchical [EFL97, KT97, STT81]. Hierarchically [LE97]. hierarchized [Car80]. Hierarchy [War77]. High [FK95a, FHH^{+}93, PT98]. implementation [KR91]. Improved [PT95a]. Incremental [Norxx, PT97, TBC97, BK97]. Influence [ELLM95]. information [Tol96]. Ink [JRZ95]. Interactive [BFGT97, DL97, Foe97, PT95b, Lew75]. InteractiveGiotto [BFGT97]. Interface [ACM90, MK91]. International [ACM89, Di 97, TT95]. Issues [PT95b]. Italy [Di 97].

Jersey [ACM89, TT95]. June [MK91].

Knowledge [NT87]. Kolmogorov [ESW96].

Labeling [KT97, KH97]. Labelings [HK95]. Large [DGK98, KRB95]. Lattices [SW95]. layered [vdP14]. layers [EW94]. Layout [Bra95, Norxx, See97, MAUY98]. Lefty [KD91, Kou94b, Kou96c]. Less [JRZ95]. Lexical [TBC97]. Line [CGKT97, EFL97, GR02, SKC96, VA88, SKC00]. Linear [Cha02, GR02, RNN97, BJL06, CP95b, Kim97, vdP14]. linear-time [CP95b]. Lines [HIIN99]. LINK [BDGS97]. Logarithmic [Kim97]. Low [FK95a]. Lower [Bie95, TTV91].

M [Rad88]. M-ary [Rad88]. Magic [Ano85]. Magnetic [SM95b]. Magnetic-Spring [SM95b]. Maintained [vOV96]. Management [ACM89]. Manipulation [BDGS97, vOV96, GS85]. Map [Saa95]. Mapping [War77]. Maps [MR95]. Method [BMT97, New89, SM95b]. Methods [Cla90, Ead96, STT81]. Microsoft [VD94]. Minimum [BDD98, CN95, CP97a, DL97, FK95b, GT97, LL95, LL97, Tre96b, Tre96a]. Minimum-Area [CP97a, Tre96b, Tre96a]. Minimum-Width [CN95]. model [KR91].

Navigation [ECH97, TBC97]. Near [Cha02]. Near-Linear [Cha02]. NEATO [Nor02b, Nor92]. Networks [TX95]. Nice [BGRT95, FK97b]. Nicely [BKW89, BKW90, SR83, ABJT04, DH96]. no [Val97]. Nodes [Wal80, Wal91]. non [vdP14]. non-layered [vdP14]. Nonplanar [FK97a]. note [Tre96b, Tre96a]. Number [BDD98, DL97]. Numbers [FK95a].

Object [See97, KS90]. Object-Oriented [See97, KS90]. Objects [BGRT95, VA88]. October [ACM89, ACM90, TT95]. octree [VA88]. octree-represented [VA88]. Online [ECH97]. open [GN00]. Optigraph [Ign95]. Optimal [Bie96, CP95c, DLV93, RNN97]. Optimal-Area [CP95c]. Optimizing [CGKT97]. Oriented [See97, KS90]. Orthogonal [BDD98, BK94, Bie95, Bie96, BSWW97, BMT97, BFGT97, CP95a, CGKT97, DLV93, DL97, EF97, ESW97, FK95b, FK97a, Foe97, PT95a, PT95b, PT97, PT98, RNN97, ESW96, TTV91]. Outerplanar [LL95].

package [VD93, VD94]. Pairwise [Val97]. Parallel [CDTT95, DLV93, HIIN99, MPS94]. Parsing [MSG95]. Passau [Bra96]. Phase [BMT97]. PhyloDraw [CJKC00]. phylogenetic [CJKC00]. Pictures [Kou94b, Kou96c, KD91]. Pitfalls [JLM97]. Placement [FR91]. Planar [CDTT95, CG95, DV96, EG95, FCE95, FK95b, FKK97, FK97b, GT94b, HK95, Hop71, LE97, CP95b, HL96, TTV91, DLV93]. Plane [Bie96, CN95, Dji95, FHH^{+}93, JRZ95, RNN96, RNN97, DT88, KH97, SSH^{+}98]. Planes [HKR95]. posets [DN97]. Positioning [Wal80, Wal91]. Power [Par95]. PQ [JLM97]. PQ-Trees [JLM97]. preprocessor [Sey92]. Pretty [Vau80]. Pretty-Printing [Vau80]. Primitives [GS85]. Princeton [ACM89, TT95]. Printing [Vau80]. Problem [Saa95]. problems [KH97]. Proceedings [ACM89, ACM90, Bra96, Di 97, MK91, NS87, Nor97, TT95]. Program [Hop71, Lan82, GNV88, Swe78]. Projections [BGRT95]. Prototyping [BDLT95]. Proximity [LD95].

Query [TBC97].

Randomized [BHR95]. Ratio [CGKT97, DGK98, GR02]. readability [TdB88]. Real [Par95]. Real-Time [Par95]. Recognizing [ELLM95]. Rectangle [ELLM95]. Rectangular [RNN96]. Refinement [TBC97]. Regular [HK95, KH97]. Relations [vOV96]. Report [EM95]. Representations [DLV93, DT88]. represented [VA88]. Requirements [KS90, LE97]. Resolution [FHH^{+}93, GT94b]. Rome [Di 97]. rooted [BJL06]. Routine [Lev90].

Screens [KRB95]. Search [CP97b]. September [Bra96, Di 97, NS87, Nor97]. Series [CDTT95, DLV93]. Series-Parallel [CDTT95, DLV93]. Service [BGT97]. SIGGRAPH [ACM90]. Simple [KRB95, SM95b]. Simulated [CT95, DH96]. Simulator [Par95]. single [HL96]. single-source [HL96]. Small [KRB95, PT98]. Snowbird [ACM90]. Software [ACM89, ACM90, NS87, See97, GN00, VD93, VD94]. source [HL96]. Space [BGRT95, LD95]. Specification [Mal82]. Spheres [HKR95]. Spirality [DLV93]. Spline [vOV96]. Splines [vOV96]. Spring [SM95b]. ST [CDTT95]. ST-Digraphs [CDTT95]. static [BK97]. Sticky [vOV96]. Straight [CGKT97, EFL97, GR02, SKC96, SKC00]. Straight-Line [CGKT97, EFL97, GR02, SKC96, SKC00]. Strasbourg [NS87]. Stressed [EG95]. structural [Fuj95]. structure [BL86].

Structures [Ber95, vOV96, STT81, Fra81, Leg91, Lew75]. Study [DGLP97]. subdivisions [GS85]. Sugiyama [See97]. supported [KR91]. Surfaces [Žit94]. Symmetrically [HE03]. Symmetry [MACL95]. Symposium [ACM90, Bra96, Di 97, Nor97]. Syntax [Fra81, Thi97]. Syntax-directed [Fra81]. System [Ber95, BDLT95, HMPT97, MACL95, MSG95, Par95, STT81, CJKC00, GN00, Lew75].

Tcl [EN96]. TclDG [EN96]. technical [KD91]. Technique [GKNV93]. techniques [ESW96]. Technology [ACM90]. Telecommunication [TX95]. Testing [BDLT95]. text [MB91]. Their [DGK98]. Theory [War77]. Third [ACM90]. Three [BMT97, CELR95, EG95, EHW97, ESW97, HE03, PTT97, PT97, PV97, ESW96]. Three-Dimensional [CELR95, EHW97, PTT97, ESW96]. Three-Phase [BMT97]. Tidier [RT81]. Tidy [Rad88, WS79, vdP14]. Time [Par95, BJL06, CP95b, vdP14]. Tool [DL97, PV97]. Topological [vOV96]. Toroidal [MR95]. Translation [BGT97]. Tree [CGKT97, HIIN99, SKC96, Til81, CJKC00, KS90, Leg91, Lew75, Pen02, SKC00, Swo91, Lew75]. tree/graph [KS90]. TREECON [VD94, VD93]. Trees [BKW89, BKW90, Cha02, CDTT95, CP95c, CP97a, CP97b, CG95, DGK98, GR02, HE03, LD95, MPS94, Moe90, Rad88, RT81, SR83, Vau80, Wal80, Wal91, WS79, ABJT04, BJL06, Kim97, MAUY98, SSH^{+}98, Tre96b, Tre96a, VD93, VD94, vdP14, JLM97, Epp85]. Triangular [DV96]. Triangulations [LL95, LL97]. Triconnected [Bie96, RNN97]. Tutorial [Nor02a]. Two [ESW97, DN97, EW94, KD91, Pen02]. two-dimensional [DN97]. two-view [KD91]. Types [Mal82].

UIST [ACM90]. UML [See97]. Understanding [STT81]. undirected [KK89]. Unified [BMT97, SM95b]. Upward [CP95c, CP97b, FK95b, HKR95, HL96, JRZ95, MR95, SKC96, Kim97, Tre96b, Tre96a]. USA [ACM90, Nor97, TT95]. Use [DGK98]. User [ACM90, Nor92]. Uses [Cla90]. Using [Ign95, JRZ95, JLM97, MSG95, TBC97, DH96]. Utah [ACM90].

v [CP97a]. Validating [PCJ95]. Very [DGK98]. view [KD91]. Viewpoints [EHW97]. Visibility [FKK97]. Visual [STT81]. Visualization [NK94, GN00, Tol96]. Volume [CS97]. Voronoi [GS85].

Web [ECH97]. Weight [LL95, LL97]. Width [CN95, Kim97]. Windows [VD94]. work [KR91]. Workbench [BDLT95]. Workshop [ACM89, TT95]. WWW [BGT97].

References

Akkerman:2004:CDT

[ABJT04] Thorsten Akkerman, Christoph Buchheim, Michael Jünger, and Daniel Teske. On the complexity of drawing trees nicely: corrigendum. Acta Informatica, 40(8):603– 607, July 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

www.springerlink.com/openurl.
asp?genre=article&issn=0001-
5903&volume=40&issue=8&spage=

603. See [SR83].

ACM:1989:PIW

[ACM89] ACM, editor. Proceedings of the 2nd International Workshop on Software Configuration Management, Princeton, New Jersey, October 24, 1989. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-334-5. LCCN QA76.76.C69 I58 1989.

ACM:1990:UTA

[ACM90] ACM, editor. UIST. Third Annual Symposium on User Interface Software and Technology. Proceedings of the ACM SIGGRAPH Symposium, Snowbird, Utah, USA, October 3–5, 1990. ACM Press, New York, NY 10036, USA, October 1990. CODEN CGRADI, CPGPBZ. ISBN 0-89791-410-4. ISSN 0097-8930. LCCN QA76.9.U83A27 1990.

URL http://info.acm.org/pubs/
contents/proceedings/graph/.

Anonymous:1985:M

[Ano85] Anonymous.  Magic.  Techni-

cal report, University of California, Berkeley, Berkeley, CA, 1985. Discusses a tool for drawing VLSI layouts.

Andalman:1997:DGB

[ARRM97] B. Andalman, K. Ryall, W. Ruml, and J. Marks. Design gallery browsers based on 2D and 3D graph drawing. Lecture Notes in Computer Science, 1353:322–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertolazzi:1998:COD

[BDD98] P. Bertolazzi, G. Di Battista, and W. Didimo. Computing orthogonal drawings with the minimum number of bends. Lecture Notes in Computer Science, 1272:331–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berry:1997:GDM

[BDGS97] J. Berry, N. Dean, M. Goldberg, and G. Shannon. Graph drawing and manipulation with LINK. Lecture Notes in Computer Science, 1353:425–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buti:1995:GSP

[BDLT95] L. Buti, G. Di Battista, G. Liotta, and E. Tassinari. GD-Workbench: a system for prototyping and testing graph drawing algorithms. Lecture Notes in Computer Science, 1027: 111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertault:1995:ADS

[Ber95] F. Bertault. Adocs: a drawing system for generic combinatorial structures. Lecture Notes in Computer Science, 1027:24–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bridgeman:1997:IAI

[BFGT97] S. S. Bridgeman, J. Fanto, A. Garg, and R. Tamassia. InteractiveGiotto: An algorithm for interactive orthogonal graph drawing. Lecture Notes in Computer Science, 1353:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1995:DNP

[BGRT95] P. Bose, F. Gomez, P. Ramos, and G. Toussaint. Drawing nice projections of objects in space. Lecture Notes in Computer Science, 1027:

52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bridgeman:1997:GDT

[BGT97] S. Bridgeman, A. Garg, and R. Tamassia. A graph drawing and translation service on the WWW. Lecture Notes in Computer Science, 1190:45–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1995:ECF

[BHR95] F. J. Brandenburg, M. Himsolt, and C. Rohrer. An experimental comparison of force-directed and randomized graph drawing algorithms. Lecture Notes in Computer Science, 1027:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1995:NLB

[Bie95] T. C. Biedl. New lower bounds for orthogonal graph drawings. Lecture Notes in Computer Science, 1027: 28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1996:OOD

[Bie96] T. C. Biedl. Optimal orthogonal drawings of triconnected plane graphs. Lecture Notes in Computer Science, 1097:333–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchheim:2006:DRT

[BJL06] Christoph Buchheim, Michael Jünger, and Sebastian Leipert. Drawing rooted trees in linear time.

Software—Practice and Experience, 36(6):651–665, May 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Biedl:1994:BHO

[BK94] T. Biedl and G. Kant. A better heuristic for orthogonal graph drawings. Lecture Notes in Computer Science, 855:24–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1997:ASI

[BK97] T. C. Biedl and M. Kaufmann.

Area-efficient static and incremental graph drawings. Lecture Notes in Computer Science, 1284:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brueggeman-Klein:1989:DTN

[BKW89] Anne Brüggeman-Klein and Derrick Wood. Drawing trees nicely with TEX. Electronic Publishing —Origination, Dissemination, and Design, 2(2):101–115, July 1989. CODEN EPODEU. ISSN 0894-3982.

Brueggeman-Klein:1990:DTN

[BKW90] Anne Brüggeman-Klein and Derrick Wood. Drawing trees nicely with TEX. In Clark [Cla90], pages 185–206. ISBN 0-13-912296-6. LCCN Z253.4.T47 C45 1990. US$69.95.

Burkhard:1986:DAS

[BL86] B. Burkhard and M. Lee. Drawing Ada structure charts. ACM SIGADA Ada Letters, 6(3):71–80, June 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Biedl:1997:TMU

[BMT97] T. C. Biedl, B. P. Madden, and I. G. Tollis. The three-phase method: a unified approach to orthogonal graph drawing. Lecture Notes in Computer Science, 1353: 391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1995:DGD

[Bra95] F. J. Brandenburg. Designing graph drawings by layout graph grammars. Lecture Notes in Computer Science, 894:416–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1996:GDS

[Bra96] F. J. (Franz J.) Brandenburg, editor. Graph drawing: Symposium on Graph Drawing, GD ’95, Passau, Germany, September 20–22, 1995: proceedings, volume 1027 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60723-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 1995.

Biedl:1997:OGD

[BSWW97] T. Biedl, T. Shermer, S. Whitesides, and S. Wismath. Orthogonal 3-D graph drawing. Lecture Notes in Computer Science, 1353:76–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpano:1980:ADH

[Car80] M. Carpano. Automatic display of hierarchized graphs for computer aided decision analysis. IEEE Transactions on Software Engineering, SE-12(4):538–546, April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Cohen:1995:DGD

[CDTT95] Robert F. Cohen, Giuseppe Di Battista, Roberto Tamassia, and Ioannis G. Tollis. Dynamic graph drawings: Trees, series-parallel digraphs, and planar ST-Digraphs. SIAM Journal on Computing, 24 (5):970–1001, ???? 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:1995:TGD

[CELR95] R. F. Cohen, P. Eades, T. Lin, and F. Ruskey. Three-dimensional graph drawing. Lecture Notes in Computer Science, 894:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cruz:1995:DGE

[CG95] I. F. Cruz and A. Garg. Drawing graphs by example efficiently: Trees and planar acyclic diagraphs. Lecture Notes in Computer Science, 894:404–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:OAA

[CGKT97] T. Chan, M. T. Goodrich, S. R.

Kosaraju, and R. Tamassia. Optimizing area and aspect ratio in straight-line orthogonal tree drawings. Lecture Notes in Computer Science, 1190:63–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:2002:NLA

[Cha02] T. M. Chan. A near-linear area bound for drawing binary trees. Algorithmica, 34(1):1–13, July 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.

springerlink.com/openurl.asp?
genre=article&issn=0178-4617&
volume=34&issue=1&spage=1.

Choi:2000:PPT

[CJKC00] J-H. Choi, H-Y. Jung, H-S. Kim, and H-G. Cho. PhyloDraw: a phylogenetic tree drawing system. Bioinformatics, 16(11):1056–1058, November 2000. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Clark:1990:TAP

[Cla90] Malcolm Clark, editor. TEX— Applications, Uses, Methods. Ellis Horwood, New York, NY, USA, 1990. ISBN 0-13-912296-6. LCCN Z253.4.T47 C45 1990. US$69.95.

Chrobak:1995:MGD

[CN95] M. Chrobak and S. Nakano. Minimum-width grid drawings of plane graphs. Lecture Notes in Computer Science, 894:104–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calamoneri:1995:EOG

[CP95a] T. Calamoneri and R. Petreschi. An efficient orthogonal grid drawing algorithm for cubic graphs. Lecture Notes in Computer Science, 959:31– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrobak:1995:LTA

[CP95b] M. Chrobak and T. H. Payne. A linear-time algorithm for drawing a planar graph on a grid. Information Processing Letters, 54(4):241–246, May 26, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crescenzi:1995:OUD

[CP95c] P. Crescenzi and A. Piperno.

Optimal-area upward drawings of AVL trees. Lecture Notes in Computer Science, 894:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1997:MHD

[CP97a] P. Crescenzi and P. Penna.

Minimum-area h-v drawings of complete binary trees. Lecture Notes in Computer Science, 1353:371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1997:UDS

[CP97b] P. Crescenzi and P. Penna. Upward drawings of search trees. Lecture Notes in Computer Science, 1197: 114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calamoneri:1997:DGO

[CS97] T. Calamoneri and A. Sterbini. Drawing 2-, 3- and 4-colorable graphs in O(n^{2}) volume. Lecture Notes in Computer Science, 1190: 53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cruz:1995:GDS

[CT95] I. F. Cruz and J. P. Twarog. 3D graph drawing with simulated annealing. Lecture Notes in Computer Science, 1027:162–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duncan:1998:BAR

[DGK98] C. A. Duncan, M. T. Goodrich, and S. G. Kobourov. Balanced aspect ratio trees and their use for drawing very large graphs. Lecture Notes in Computer Science, 1547:111–124, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBattista:1997:DDA

[DGLP97] G. Di Battista, A. Garg, G. Liotta, and A. Parise. Drawing directed acyclic graphs: An experimental study. Lecture Notes in Computer Science, 1190:76–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davidson:1996:DGN

[DH96] Ron Davidson and David Harel. Drawing graphs nicely using simulated annealing. ACM Transactions on Graphics, 15(4):301–331, October 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0730-
0301/234538.html.

DiBattista:1997:GDI

[Di 97] Giuseppe Di Battista, editor. Graph drawing: 5th International Symposium, GD’97, Rome, Italy, September 18–20, 1997: proceedings, volume 1353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63938-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1353.

Djidjev:1995:DGC

[Dji95] H. N. Djidjev. On drawing a graph convexly in the plane. Lecture Notes in Computer Science, 894:76– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Didimo:1997:GIT

[DL97] W. Didimo and A. Leonforte.

GRID: An interactive tool for computing orthogonal drawings with the minimum number of bends. Lecture Notes in Computer Science, 1353:309–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBattista:1993:SOR

[DLV93] G. Di Battista, G. Liotta, and F. Vargiu. Spirality of orthogonal representations and optimal drawings of series-parallel graphs and 3-Planar graphs. Lecture Notes in Computer Science, 709:151–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLaHiguera:1997:DET

[DN97] C. De La Higuera and L. Nourine. Drawing and encoding two-dimensional posets. Theoretical Computer Science, 175(2):293–308, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DiBattista:1988:APR

[DT88] Giuseppe Di Battista and Roberto Tamassia. Algorithms for plane representations of acyclic digraphs. Theoretical Computer Science, 61(2-3):175–198, November 1988. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DiBattista:1996:APT

[DV96] Giuseppe Di Battista and Luca Vismara. Angles of planar triangular graphs. SIAM Journal on Discrete Mathematics, 9(3):349–359, August 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Eades:1984:HGD

[Ead84] Peter Eades. A heuristic for graph drawing. Congressus Numerantium, 42(??):149–160, ???? 1984. ISSN 0384-9864.

Eades:1996:GDM

[Ead96] P. Eades. Graph drawing methods. Lecture Notes in Computer Science, 1115:40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:OAG

[ECH97] P. Eades, R. F. Cohen, and Mao Lin Huang. Online animated graph drawing for Web navigation. Lecture Notes in Computer Science, 1353: 330–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:DCG

[EF97] P. Eades and Q.-W Feng. Drawing clustered graphs on an orthogonal grid. Lecture Notes in Computer Science, 1353:146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:SDA

[EFL97] P. Eades, Q. W. Feng, and X. Lin.

Straight-line drawing algorithms for hierarchical graphs and clustered graphs. Lecture Notes in Computer Science, 1190:113–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1995:DSP

[EG95] P. Eades and P. Garvan. Drawing stressed planar graphs in three dimensions. Lecture Notes in Computer Science, 1027:212–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:FBV

[EHW97] P. Eades, M. E. Houle, and R. Webber. Finding the best viewpoints for three-dimensional graph drawings. Lecture Notes in Computer Science, 1353:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ElGindy:1995:RRI

[ELLM95] H. ElGindy, G. Liotta, A. Lubiw, and H. Meijer. Recognizing rectangle of influence drawable graphs. Lecture Notes in Computer Science, 894:352–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1995:GDC

[EM95] P. Eades and J. Marks. Graph drawing contest report. Lecture Notes in Computer Science, 894:143–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellson:1996:TTE

[EN96] John Ellson and Stephen North. TclDG — a Tcl extension for dynamic graphs. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1996. 11 pp. URL

                                        http://www.research.att.com/
                                        sw/tools/graphviz/EN96.pdf.

Eppstein:1985:TT

[Epp85] David Eppstein. Trees in TEX. TUGboat, 6(1):31–35, March 1985. ISSN 0896-3207.

Eades:1996:TKB

[ESW96] Peter Eades, Charles Stirk, and Sue Whitesides. The techniques of Kolmogorov and Barzdin for threedimensional orthogonal graph drawings. Information Processing Letters, 60(2):97–103, ???? 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eades:1997:TAT

[ESW97] P. Eades, A. Symvonis, and S. Whitesides. Two algorithms for three dimensional orthogonal graph drawing. Lecture Notes in Computer Science, 1190:139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1994:DGT

[EW94] P. Eades and S. Whitesides. Drawing graphs in two layers. Theoretical Computer Science, 131(2):361– 374, September 12, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feng:1995:HDP

[FCE95] Q.-W. Feng, R. F. Cohen, and P. Eades. How to draw a planar clustered graph. Lecture Notes in Computer Science, 959:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formann:1993:DGP

[FHH^{+}93] M. Formann, T. Hagerup, J. Haralambides, M. Kaufmann, F. T. Leighton, A. Symvonis, E. Welzl, and G. Woeginger. Drawing graphs in the plane with high resolution. SIAM Journal on Computing, 22(5): 1035–1052, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Foessmeier:1995:DHD

[FK95a] U. Foessmeier and M. Kaufmann.

Drawing high degree graphs with low bend numbers. Lecture Notes in Computer Science, 1027:254–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1995:BOU

[FK95b] U. Foessmeier and M. Kaufmann. On bend-minimum orthogonal upward drawing of directed planar graphs. Lecture Notes in Computer Science, 894:52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:AAB

[FK97a] U. Foessmeier and M. Kaufmann. Algorithms and area bounds for nonplanar orthogonal drawings. Lecture Notes in Computer Science, 1353:134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:NDP

[FK97b] U. Foessmeier and M. Kaufmann. Nice drawings for planar bipartite graphs. Lecture Notes in Computer Science, 1203:122–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:DPG

[FKK97] U. Foessmeier, G. Kant, and M. Kaufmann. 2-Visibility drawings of planar graphs. Lecture Notes in Computer Science, 1190:155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:IOG

[Foe97] U. Foessmeier. Interactive orthogonal graph drawing: Algorithms and bounds. Lecture Notes in Computer Science, 1353:111–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fruchterman:1991:GDF

[FR91] Thomas M. J. Fruchterman and Edward M. Reingold. Graph drawing by force-directed placement. Software—Practice and Experience, 21(11):1129–1164, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fraser:1981:SDE

[Fra81] C. W. Fraser. Syntax-directed editing of general data structures. ACM SIGPLAN Notices, 16(6):17– 21, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fujita:1995:XDC

[Fuj95] Shinsaku Fujita. XΥMTEX for drawing chemical structural formulas. TUGboat, 16(1):80–88, March 1995. ISSN 0896-3207.

Gansner:2003:DGG

[Gan03] Emden R. Gansner. Drawing graphs with GraphViz. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 15, 2003. URL

http://www.research.att.com/
sw/tools/graphviz/libguide.pdf.

Garg:1995:DAG

[Gar95] A. Garg. On drawing angle graphs.

Lecture Notes in Computer Science, 894:84–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gansner:2002:DGD

[GKN02] Emden Gansner, Eleftherios Koutsofios, and Stephen North. Drawing graphs with dot. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 4, 2002. 40 pp. URL

http://www.research.att.com/
sw/tools/graphviz/dotguide.pdf.

Gansner:1993:TDD

[GKNV93] Emden R. Gansner, Eleftherios Koutsofios, Stephen C. North, and Kiem-Phong Vo. A technique for drawing directed graphs. IEEE Transactions on Software Engineering, 19(3):214–230, May 1993. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://www.research.att.
com/sw/tools/graphviz/TSE93.
pdf.

Gansner:2000:OGV

[GN00] Emden R. Gansner and Stephen C. North. An open graph visualization system and its applications to software engineering. Software—Practice and Experience, 30(11):1203– 1233, September 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www.research.att.com/
sw/tools/graphviz/GN99.pdf;
http://www3.interscience.wiley.
com/cgi-bin/abstract/72517075/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=
72517075&PLACEBO=IE.pdf.

Gansner:1988:DPD

[GNV88] E. R. Gansner, S. C. North, and K. P. Vo. DAG: a program that draws directed graphs. Software —Practice and Experience, 18(11): 1047–1062, November 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Garg:2002:SLD

[GR02] Ashim Garg and Adrian Rusu.

Straight-line drawings of binary trees with linear area and arbitrary aspect ratio. Lecture Notes in Computer Science, 2528:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.de/
link/service/series/0558/bibs/
2528/25280320.htm; http://link.
springer.de/link/service/series/
0558/papers/2528/25280320.pdf.

Gropp:1995:DC

[Gro95] H. Gropp. The drawing of configurations. Lecture Notes in Computer Science, 1027:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guibas:1985:PMG

[GS85] Leonidas Guibas and Jorge Stolfi.

Primitives for the manipulation of general subdivisions and computation of Voronoi diagrams. ACM Transactions on Graphics, 4(2):74– 123, April 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garg:1994:AGD

[GT94a] A. Garg and R. Tamassia. Advances in graph drawing. Lecture Notes in Computer Science, 778:12–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1994:PDA

[GT94b] A. Garg and R. Tamassia. Planar drawings and angular resolution: Algorithms and bounds. Lecture Notes in Computer Science, 855:12–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1997:NMC

[GT97] A. Garg and R. Tamassia. A new minimum cost flow algorithm with applications to graph drawing. Lecture Notes in Computer Science, 1190:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:2003:DTS

[HE03] Seok-Hee Hong and Peter Eades. Drawing trees symmetrically in three dimensions. Algorithmica, 36(2):153–178, April 2003. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.asp?
                                        genre=article&issn=0178-4617&
                                        volume=36&issue=2&spage=153.

Hirabayashi:1999:DTP

[HIIN99] R. Hirabayashi, Y. T. Ikebe, K. Iwamura, and T. Nakayama. Drawing a tree on parallel lines. Computers and Mathematics with Applications, 37(11):171–176, June 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Himsolt:199x:G

[Him9x] M. Himsolt. Graphed 3.0, 199x. Available by anonymous ftp at ftp://forwiss.unipassau.de/pub/local/graphed.

He:1995:REL

[HK95] X. He and M.-Y. Kao. Regular edge labelings and drawings of planar graphs. Lecture Notes in Computer Science, 894:96–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hashemi:1995:UDP

[HKR95] S. M. Hashemi, A. Kisielewicz, and I. Rival. Upward drawings on planes and spheres. Lecture Notes in Computer Science, 1027:277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutton:1996:UPD

[HL96] Michael D. Hutton and Anna Lubiw. Upward planar drawing of single-source acyclic digraphs. SIAM Journal on Computing, 25(2): 291–311, April 1996. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hundack:1997:AGD

[HMPT97] C. Hundack, P. Mutzel, I. Pouchkarev, and S. Thome. ArchE: a graph drawing system for archaeology. Lecture Notes in Computer Science, 1353:297–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hope:1971:PGD

[Hop71] Alexander K. Hope. A planar graph drawing program. Software —Practice and Experience, 1(1):83– 91, January/March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ignatowicz:1995:DFG

[Ign95] J. Ignatowicz. Drawing forcedirected graphs using Optigraph. Lecture Notes in Computer Science, 1027:333–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juenger:1997:PUP

[JLM97] M. Juenger, S. Leipert, and P. Mutzel. Pitfalls of using PQ-Trees in automatic graph drawing. Lecture Notes in Computer Science, 1353:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourdan:1995:UDP

[JRZ95] G.-V. Jourdan, I. Rival, and N. Zaguia. Upward drawing on the plane grid using less ink. Lecture Notes in Computer Science, 894:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koutsofios:1991:LTV

[KD91] Eleftherios Koutsofios and David Dobkin. Lefty: a two-view editor for technical pictures. In MacKay and Kidd [MK91], pages 67–76. ISBN 0-9695338-0-2. LCCN T385.P792 1991.

Kant:1997:REL

[KH97] G. Kant and Xin He. Regular edge labeling of 4-connected plane graphs and its applications in graph drawing problems. Theoretical Computer Science, 172(1-2):175–193, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:1997:LWL

[Kim97] Sung Kwon Kim. Logarithmic width, linear area upward drawing of AVL trees. Information Processing Letters, 63(6):303–307, ???? 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kamada:1989:ADG

[KK89] Tomihisa Kamada and Satoru Kawai. An algorithm for drawing general undirected graphs. Information Processing Letters, 31(1):7–15, April 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koutsofios:1991:DGD

[KN91] Eleftherios Koutsofios and Stephen North. Drawing graphs with dot. Technical Report 910904-59113-08TM, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1991.

Koutsofios:1994:EGD

[Kou94a] Eleftherios Koutsofios. Editing graphs with dotty. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 5, 1994. This report, and the program, is included in the graphviz package, available for non-commercial

use at http://www.research.att.
com/sw/tools/graphviz/.

Koutsofios:1994:EPL

[Kou94b] Eleftherios Koutsofios. Editing pictures with lefty. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 5, 1994. This report, and the program, is included in the graphviz package, available for non-commercial

use at http://www.research.att.
com/sw/tools/graphviz/.

Koutsofios:1996:DGD

[Kou96a] Eleftherios Koutsofios. Drawing graphs with dot. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 20, 1996. This report, and the program, is included in the graphviz package, available for non-commercial

use at http://www.research.att.
com/sw/tools/graphviz/.

Koutsofios:1996:EGD

[Kou96b] Eleftherios Koutsofios. Editing graphs with dotty. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 24, 1996. 32 pp. URL

http://www.research.att.com/
sw/tools/graphviz/dottyguide.
pdf.

Koutsofios:1996:EPL

[Kou96c] Eleftherios Koutsofios. Editing pictures with lefty. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 24, 1996. 43 pp. URL

http://www.research.att.com/
sw/tools/graphviz/leftyguide.
pdf.

Krishnamurthy:1991:EAM

[KR91] Balachander Krishnamurthy and David S. Rosenblum. An eventaction model of computer-supported cooperative work: Design and implementation. In K. Gorling and C. Sattler, editors, Proceedings of the International Workshop on Computer Supported Cooperative Work, pages 132–145. ????, ????, 1991. IFIP TC 6/WG C.5.

Kaugars:1995:SAD

[KRB95] K. Kaugars, J. Reinfelds, and A. Brazma. A simple algorithm for drawing large graphs on small screens. Lecture Notes in Computer Science, 894:278–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karrer:1990:REO

[KS90] A. Karrer and W. Scacchi. Requirements for an extensible objectoriented tree/graph editor. In ACM [ACM90], pages 84–91. CODEN CGRADI, CPGPBZ. ISBN 0-89791-410-4. ISSN 0097-8930. LCCN QA76.9.U83A27 1990.

URL http://info.acm.org/pubs/
contents/proceedings/graph/.

Kakoulis:1997:ALE

[KT97] K. C. Kakoulis and I. G. Tollis. An algorithm for labeling edges of hierarchical drawings. Lecture Notes in Computer Science, 1353:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1982:APD

[Lan82] T. Lang. Another program for drawing diagrams. Software—Practice and Experience, 12(5):397–404, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liotta:1995:CPD

[LD95] G. Liotta and G. Di Battista. Computing proximity drawings of trees in the 3-Dimensional space. Lecture Notes in Computer Science, 955: 239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1997:ARD

[LE97] X. Lin and P. Eades. Area requirements for drawing hierarchically planar graphs. Lecture Notes in Computer Science, 1353:219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leguy:1991:DTS

[Leg91] Bernard Leguy. Drawing tree structures with GWEZ. Cahiers GUTenberg, 10–11:135–146, September 1991. ISSN 1140-9304.

Levine:1990:LGR

[Lev90] Michael J. S. Levine. A LaTEX graphics routine for drawing Feynman diagrams. Computer Physics Communications, 58(1 and 2):181– 198, February/March 1990. Available in the CPC library, and

                                        from listral@ib.rl.ac.uk with

an e-mail request GET CPC INTRO

                                        CPCINDEX.

Lewis:1975:TIS

[Lew75] J. W. Lewis. TREE: An interactive system for editing tree structures. Computers and Graphics, 1(1):65– 68, May 1975. CODEN COGRD2. ISSN 0097-8493.

Lenhart:1995:HDO

[LL95] W. Lenhart and G. Liotta. How to draw outerplanar minimum weight triangulations. Lecture Notes in Computer Science, 1027:373–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenhart:1997:DFM

[LL97] W. Lenhart and G. Liotta. Drawable and forbidden minimum weight triangulations. Lecture Notes in Computer Science, 1353:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manning:1995:SDG

[MACL95] J. Manning, M. Atallah, K. Cudjoe, and J. Lozito. A system for drawing graphs with geometric symmetry. Lecture Notes in Computer Science, 894:262–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mallgren:1982:FSG

[Mal82] William R. Mallgren. Formal specification of graphic data types. ACM Transactions on Programming Languages and Systems, 4(4):687–710, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Miyadera:1998:DFL

[MAUY98] Youzou Miyadera, Koushi Anzai, Hiroshi Unno, and Takeo Yaku. Depth-first layout algorithm for trees. Information Processing Letters, 66(4):187–194, May 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maclenan:1991:ADC

[MB91] M. P. Maclenan and G. M. Burns. An approach to drawing circuit diagrams for text books. TUGboat, 12 (1):66–69, March 1991. ISSN 0896-3207.

MacKay:1991:GIC

[MK91] S. MacKay and E. M. Kidd, editors. Graphics Interface ’91, Calgary, Alberta, 3–7 June 1991: proceedings. Canadian Information Processing Society, 243 College St, 5th Floor, Toronto, Ontario M5T 2Y1, Canada, 1991. ISBN 0-9695338-0-2. LCCN T385.P792 1991.

Moen:1990:DDT

[Moe90] Sven Moen. Drawing dynamic trees. IEEE Software, 7(4):21–28, July 1, 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Metaxas:1994:PHD

[MPS94] P. T. Metaxas, G. E. Pantziou, and A. Symvonis. Parallel h-v drawings of binary trees. Lecture Notes in Computer Science, 834:487–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohar:1995:FAU

[MR95] B. Mohar and P. Rosenstiehl. A flow approach to upward drawings of toroidal maps. Lecture Notes in Computer Science, 894:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCreary:1995:CGD

[MSG95] C. McCreary, F.-S. Shieh, and H. Gill. CG: a graph drawing system using graph-grammar parsing. Lecture Notes in Computer Science, 894:270–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Newbery:1989:ECM

[New89] Frances J. Newbery. Edge concentration: a method for clustering directed graphs. In ACM [ACM89], pages 78–85. ISBN 0-89791-334-5. LCCN QA76.76.C69 I58 1989. Also published as ACM SIGSOFT Software Engineering Notes, vol. 17, no. 7, November 1989.

North:1994:AGV

[NK94] Stephen C. North and Eleftherios Koutsofios. Applications of graph visualization. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1994. URL

http://www.research.att.com/
sw/tools/graphviz/GI94.pdf.

North:1992:NUG

[Nor92] Stephen C. North. NEATO user’s guide. Technical Report 59113-921014-14TM, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 14, 1992. This report, and the program, is included in the graphviz package, available for non-commercial

use at http://www.research.att.
com/sw/tools/graphviz/.

North:1997:GDS

[Nor97] Stephen C. North, editor. Graph drawing: Symposium on Graph Drawing, GD ’96, Berkeley, California, USA, September 18–20, 1996: proceedings, volume 1190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62495-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .S884 1996.

North:19xx:ILD

[Norxx] Stephen C. North. Incremental layout in DynaDAG. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, 19xx. URL

http://www.research.att.com/
sw/tools/graphviz/dynadag.pdf.

North:2002:AT

[Nor02a] Stephen C. North. Agraph tutorial. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 2, 2002. 22 pp.

URL http://www.research.att.
com/sw/tools/graphviz/Agraph.
pdf.

North:2002:DGN

[Nor02b] Stephen C. North. Drawing graphs with NEATO. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 10, 2002. 11 pp. URL

http://www.research.att.com/
sw/tools/graphviz/neatoguide.
pdf.

Nichols:1987:EES

[NS87] H. K. (Howard K.) Nichols and D. (Dan) Simpson, editors. ESEC ’87: 1st European Software Engineering Conference, Strasbourg, France, September 9–11, 1987: proceedings. Springer-Verlag Inc., New York, NY, USA, 1987. ISBN 0-387-18712-X. LCCN QA76.758 .E97 1987.

Newbery:1987:KBE

[NT87] Frances Newbery and Walter Tichy.

Knowledge based editors for directed graphs. In Nichols and Simpson [NS87], pages 101–109. ISBN 0-387-18712-X. LCCN QA76.758 .E97 1987.

Ohl:1995:DFD

[Ohl95] Thorsten Ohl. Drawing Feynman diagrams with LaTEX and - . Computer Physics Communications, 90(2–3):340–354, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/001046559590137S.

Paris:1995:ADC

[Par95] G. Paris. Automatic drawing of compound digraphs for a real-time power system simulator. Lecture Notes in Computer Science, 1027: 431–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Purchase:1995:VGD

[PCJ95] H. C. Purchase, R. F. Cohen, and M. James. Validating graph drawing aesthetics. Lecture Notes in Computer Science, 1027:435–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Penna:2002:ATT

[Pen02] P. Penna. On the approximability of two tree drawing conventions. Information Processing Letters, 82(5): 237–242, June 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Papakostas:1995:IAB

[PT95a] A. Papakostas and I. G. Tollis. Improved algorithms and bounds for orthogonal drawings. Lecture Notes in Computer Science, 894:40– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1995:IIO

[PT95b] A. Papakostas and I. G. Tollis. Issues in interactive orthogonal graph drawing. Lecture Notes in Computer Science, 1027:419–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1997:IOG

[PT97] A. Papakostas and I. G. Tollis. Incremental orthogonal graph drawing in three dimensions. Lecture Notes in Computer Science, 1353:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1998:ODH

[PT98] A. Papakostas and I. G. Tollis. Orthogonal drawing of high degree graphs with small area and few bends. Lecture Notes in Computer Science, 1272:354–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pach:1997:TGD

[PTT97] J. Pach, T. Thiele, and G. Toth.

Three-dimensional grid drawings of graphs. Lecture Notes in Computer Science, 1353:47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patrignani:1997:TTD

[PV97] M. Patrignani and F. Vargiu.

3DCube: a tool for three dimensional graph drawing. Lecture Notes in Computer Science, 1353:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radack:1988:TDM

[Rad88] G. M. Radack. Tidy drawing of Mary trees. Technical Report CES-88-24, Department of Computer Engineering and Science, Case Western Reserve University, Cleveland, OH, USA, November 1988.

Rowe:1987:BDG

[RDM^{+}87] Lawrence A. Rowe, Michael Davis, Eli Messinger, Carl Mayer, Charles Spirakis, and Allen Tuan. A browser for directed graphs. Software—Practice and Experience, 17 (1):61–76, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rahman:1996:RGD

[RNN96] S. M. Rahman, S.-I. Nakano, and T. Nishizeki. Rectangular grid drawings of plane graphs. Lecture Notes in Computer Science, 1090: 92–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahman:1997:LAO

[RNN97] M. Saidur Rahman, S.-I. Nakano, and T. Nishizeki. A linear algorithm for optimal orthogonal drawings of triconnected cubic plane graphs. Lecture Notes in Computer Science, 1353:99–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reingold:1981:TDT

[RT81] E. M. Reingold and J. S. Tilford. Tidier drawing of trees. IEEE Transactions on Software Engineering, SE-7(2):223–228, March 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Saalfeld:1995:MGG

[Saa95] A. Saalfeld. Map generalization as a graph drawing problem. Lecture Notes in Computer Science, 894: 444–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seemann:1997:ESA

[See97] J. Seemann. Extending the sugiyama algorithm for drawing UML class diagrams: Towards automatic layout of object-oriented software diagrams. Lecture Notes in Computer Science, 1353:415–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seyfarth:1992:DDP

[Sey92] Ray Seyfarth. Diag: a drawing preprocessor for LaTEX. TUGboat, 13 (4):478–485, December 1992. ISSN 0896-3207.

Shin:1996:AAU

[SKC96] C.-S. Shin, Sung Kwon Kim, and K.-Y. Chwa. Area-efficient algorithms for upward straight-line tree drawings. Lecture Notes in Computer Science, 1090:106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shin:2000:AEA

[SKC00] C.-S. Shin, S. K. Kim, and K.-Y.

Chwa. Area-efficient algorithms for straight-line tree drawings. Computational Geometry. Theory and Applications, 15(4):175–202, April 2000. CODEN ???? ISSN 0925-7721.

Shieh:1995:DGD

[SM95a] F.-S. Shieh and C. L. McCreary.

Directed graphs drawing by clanbased decomposition. Lecture Notes in Computer Science, 1027:472–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugiyama:1995:SUM

[SM95b] K. Sugiyama and K. Misue. A simple and unified method for drawing graphs: Magnetic-spring algorithm. Lecture Notes in Computer Science, 894:364–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Supowit:1983:CDT

[SR83] Kenneth J. Supowit and Edward M. Reingold. The complexity of drawing trees nicely. Acta Informatica, 18(4):377–392, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See corrigendum [ABJT04].

Sassani:1998:ADB

[SSH^{+}98] D. C. Sassani, E. L. Shock, K. Harstad, J. Bellan, C.-S. Shin, K. K. Sung, S.-H. Kim, and K.-Y. Chwa. Algorithms for drawing binary trees in the plane. Information Processing Letters, 66(3):133–139, May 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sugiyama:1981:MVU

[STT81] K. Sugiyama, S. Tagawa, and M. Toda. Methods for visual understanding of hierarchical system structures. IEEE Transactions on Systems, Man, and Cybernetics, SMC-11(2):109–125, February 1981. CODEN ISYMAW. ISSN 0018-9472, 1083-4427.

Stumme:1995:GHD

[SW95] G. Stumme and R. Wille. A geometrical heuristic for drawing concept lattices. Lecture Notes in Computer Science, 894:452–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sweet:1978:EEP

[Swe78] R. E. Sweet. Empirical estimates of program entropy. Technical Report STAN-CS-78-698, Department of Computer Science, Stanford University, Stanford, CA, USA, November 1978. Also issued as Report CSL-78-3, Xerox PARC, Palo Alto, CA, September 1978.

Swonk:1991:LTD

[Swo91] Glenn L. Swonk. LaTEX tree drawer. TUGboat, 12(2):286–289, June 1991. ISSN 0896-3207.

Tunkelang:1997:LNU

[TBC97] D. Tunkelang, R. J. Byrd, and J. W.

Cooper. Lexical navigation: Using incremental graph drawing for query refinement. Lecture Notes in Computer Science, 1353:316–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tamassia:1988:AGD

[TdB88] R. Tamassia, G. di Battista, and C. Batini. Automatic graph drawing and readability of diagrams. IEEE Transactions on Systems, Man, and Cybernetics, 18(1):61–79, January/ February 1988. CODEN ISYMAW. ISSN 0018-9472, 1083-4427.

Thimbleby:1996:ECA

[Thi96] Harold Thimbleby. An equivalence class algorithm for drawing autostereograms. Software—Practice and Experience, 26(3):309–325, March 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Thiemann:1997:DSD

[Thi97] P. J. Thiemann. Drawing syntax diagrams in Haskell. Software—Practice and Experience, 27(8):967–??, August 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tilford:1981:TDA

[Til81] J. S. Tilford. Tree drawing algorithms. Master’s thesis, Department of Computer Science, University of Illinois, Urbana, IL, USA, April 1981. Issued as report UIUCDCSR-81-1055, and as document UILU-ENG-81-1711 from College of Engineering Document Center at the University of Illinois.

Tollis:1996:GDI

[Tol96] Ioannis G. Tollis. Graph drawing and information visualization. ACM Computing Surveys, 28(4es): 19, December 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http:

                                        //www.acm.org/pubs/citations/
                                        journals/surveys/1996-28-4es/
                                        a19-tollis/.

Trevisan:1996:NMU

[Tre96a] L. Trevisan. A note on minimumarea upward drawing of complete and Fibonacci trees. Information Processing Letters, 57(5):231– ??, March 11, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Trevisan:1996:NMA

[Tre96b] Luca Trevisan. A note on minimumarea upward drawing of complete and Fibonacci trees. Information Processing Letters, 57(5):231–236, ???? 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tamassia:1995:GDD

[TT95] Roberto Tamassia and Ioannis G.

Tollis, editors. Graph drawing: DIMACS International Workshop, GD ’94, Princeton, New Jersey, USA, October 10–12, 1994: proceedings, volume 894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-58950-3 (Berlin), 0-387-58950-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .D55 1994. DM94.00.

Tamassia:1991:LBP

[TTV91] Roberto Tamassia, Ioannis G. Tollis, and Jeffrey Scott Vitter. Lower bounds for planar orthogonal drawings of graphs. Information Processing Letters, 39(1):35–40, July 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tollis:1995:DTN

[TX95] I. G. Tollis and C. Xia. Drawing telecommunication networks. Lecture Notes in Computer Science, 894:206–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veenstra:1988:LDO

[VA88] Jack Veenstra and Narendra Ahuja.

Line drawings of octree-represented objects. ACM Transactions on Graphics, 7(1):61–75, January 1988.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.acm.
                                        org/pubs/toc/Abstracts/0730-
                                        0301/42189.html.

Valtr:1997:GDN

[Val97] P. Valtr. Graph drawings with no k pairwise crossing edges. Lecture Notes in Computer Science, 1353: 205–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaucher:1980:PPT

[Vau80] Jean G. Vaucher. Pretty-printing of trees. Software—Practice and Experience, 10(7):553–561, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VandePeer:1993:TSP

[VD93] Y. Van de Peer and R. De Wachter. Treecon: a software package for the construction and drawing of evolutionary trees. Computer Applications in the Biosciences: CABIOS, 9(2):177–??, April 1, 1993. CODEN COABER. ISSN 0266-7061.

VandePeer:1994:TWS

[VD94] Y. Van de Peer and R. De Wachter. TREECON for Windows: a software package for the construction and drawing of evolutionary trees for the Microsoft Windows environment. Computer Applications in the Biosciences: CABIOS, 10(5):569– ??, September 1, 1994. CODEN COABER. ISSN 0266-7061.

vanderPloeg:2014:DNL

[vdP14] Atze van der Ploeg. Drawing nonlayered tidy trees in linear time.

Software—Practice and Experience, 44(12):1467–1484, December 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Overveld:1996:SSD

[vOV96] C. W. A. M. van Overveld and Marie Luce Viaud. Sticky splines: Definition and manipulation of spline structures with maintained topological relations. ACM Transactions on Graphics, 15(1):72–98, January 1996. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0730-
0301/225889.html;   http://www.
acm.org/pubs/toc/Abstracts/0730-
0301/226154.html.

Walker:1980:PNG

[Wal80] John Q. Walker II. Positioning nodes for general trees. Software— Practice and Experience, 10(7):553– 561, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Walker:1991:PNG

[Wal91] John Q. Walker II. Positioning nodes for general trees. C Users Journal, 9(2):47–62, February 1991. ISSN 0898-9788.

Warfield:1977:CTH

[War77] John Warfield. Crossing theory and hierarchy mapping. IEEE Transactions on Systems, Man, and Cybernetics, SMC-7(7):505–523, July 1977. CODEN ISYMAW. ISSN 0018-9472, 1083-4427.

Wetherell:1979:TDT

[WS79] C. S. Wetherell and A. Shannon. Tidy drawing of trees. IEEE Transactions on Software Engineering, SE-5(5):514–520, September 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Zitnik:1994:DGS

[Žit94] Arjana Žitnik. Drawing graphs on surfaces. SIAM Journal on Discrete Mathematics, 7(4):593–597, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zinssmeister:1996:DGA

[ZM96] G. Zinssmeister and C. L. Mc-Creary. Drawing graphs with attribute graph grammars. Lecture Notes in Computer Science, 1073: 443–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).